OpenStudy (anonymous):

what is the hypotenuse of a right triangle if one leg is 10 and the other leg is 13. use pythagreom therom to solve

OpenStudy (amistre64):

and what is the pythag thrm tell us?

OpenStudy (anonymous):

a2 + b2 = c2

OpenStudy (amistre64):

good, then lets fill in our numbers 10^2 + 13^2 = c^2

OpenStudy (anonymous):

wow thats is right

OpenStudy (amistre64):

the only issue might be in going from c^2 to c for that we just take the square root of each side sqrt(10^2+13^2) = c

OpenStudy (anonymous):

no that is wrong you square the 10 and the 13 and then you get 100+169 then you add all that and you get 296 and then take the square root of that and you get 16.4 for the hypotenuse you weirdo.